# Hybrid Beamforming Design for RSMA-assisted mmWave Integrated Sensing and Communications

### v1

@@ II-C Radar Sensing Model And Metrics
In ISAC systems, the radar sensing waveform multiplexes the communication signal $\boldsymbol{\rm x}$. According to [26], we propose the HBF architecture for radar sensing that combines the advantages of the high directional gain of phased-array radars and the diversity gain provided by multiple waveforms of conventional MIMO radars. The purpose of radar sensing is to realize functions such as target detection and parameter estimation in a cluttered environment. In this paper, we assume that the sensing function of the ISAC system is to detect a radar target (located at angles $\theta_{0}$) of interest in the presence of interference from $Q$ clutter located at different angles $\theta_{q},q\in\mathcal{Q}=\{1,\cdots,Q\}$. Thus, the echo signal arriving at the ISAC receiver, denoted by $\boldsymbol{\rm y}_{r}$, can be expressed as
$\boldsymbol{\rm y}_{r}=$ $\underbrace{\sqrt{N_{r}N_{t}}\xi_{0}\boldsymbol{\rm a}_{r}\left(\theta_{0}\right)\boldsymbol{\rm a}_{t}^{H}\left(\theta_{0}\right)\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm s}}_{\text{target }}+$ (8)
@@
where $\mathbf{a}_{t}(\theta)\in\mathbb{C}^{N_{{t}}\times 1}$ and $\mathbf{a}_{r}(\theta)\in\mathbb{C}^{N_{{r}}\times 1}$ denote the transmit and receive steering vectors, respectively, similar to the definition of Eq.(4). $\mathbf{z}_{r}\sim\mathcal{CN}\left(\mathbf{0}_{N_{r}},\sigma_{{z}}^{2}\mathbf{I}_{N_{r}}\right)$ represents the AWGN with zero mean and variance $\sigma_{{z}}^{2}$. $\xi_{0}$ and $\xi_{q}$ denote the complex fading coefficients of the target and the $q$-th clutter, respectively. Based on the radar equations [39], the modulus of $\xi_{0}$, denoted by $\left|\xi_{0}\right|$, can be expressed by
$\left|\xi_{0}\right|=\sqrt{\frac{G_{t}G_{r}\sigma_{0}^{RCS}\lambda^{2}}{(4\pi)^{3}R_{0}^{4}}},$ (9)
where $G_{t}$ and $G_{r}$ denote the transmit and receive antenna gains, respectively. $\sigma_{0}^{RCS}$ represents the radar cross section (RCS) of the target, and $R_{0}$ represents the distance between the target and the ISAC BS. Also, the modulus of $\xi_{q},q\in\mathcal{Q}$ is similar to Eq.(9).
@@
where $\boldsymbol{\rm A}\left(\theta_{0}\right)=\sqrt{N_{r}N_{t}}\boldsymbol{\rm a}_{r}\left(\theta_{0}\right)\boldsymbol{\rm a}_{t}^{H}\left(\theta_{0}\right)$, $\boldsymbol{\rm A}\left(\theta_{q}\right)=\sqrt{N_{r}N_{t}}\boldsymbol{\rm a}_{r}\left(\theta_{q}\right)\boldsymbol{\rm a}_{t}^{H}\left(\theta_{q}\right)$. Therefore, the filtered output SCNR, denoted by $\gamma_{r}^{out}$, can be expressed as [18],[40]:
$\gamma_{r}^{out}=$ $\frac{\left|\xi_{0}\boldsymbol{\rm v}^{H}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm x}\right|^{2}}{\mathbb{E}\left\{\left|\boldsymbol{\rm v}^{H}\sum_{q\in\mathcal{Q}}\xi_{q}\boldsymbol{\rm A}\left(\theta_{q}\right)\boldsymbol{\rm x}\right|^{2}+\left|\boldsymbol{\rm v}^{H}\boldsymbol{\rm z}_{r}\right|^{2}\right\}}$
$=$ $\frac{\left|\xi_{0}\boldsymbol{\rm v}^{H}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm x}\right|^{2}}{\boldsymbol{\rm v}^{H}\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)\boldsymbol{\rm v}},$ (11)
where $\boldsymbol{\rm R}_{c}=\sum_{q\in\mathcal{Q}}\left|\xi_{q}\right|^{2}\boldsymbol{\rm A}\left(\theta_{q}\right)\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm F}_{\rm D}^{H}\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm A}^{H}\left(\theta_{q}\right)$.The receive beamforming problem of solving the output SCNR maximization is well known as the minimum variance distortionless response (MVDR) beamforming problem [41], which has a closed-form solution, denoted by $\boldsymbol{\rm v}^{\star}$, and can be expressed as
$\boldsymbol{\rm v}^{\star}=\frac{\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)^{-1}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm x}}{\boldsymbol{\rm x}^{H}\boldsymbol{\rm A}^{H}\left(\theta_{0}\right)\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)^{-1}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm x}}.$ (12)
By substituting the optimal receive beamforming vector $\boldsymbol{\rm v}^{\star}$ into Eq.(11) and subsequently taking the expectation, we can derive the average output SCNR, denoted as $\overline{\gamma}_{r}^{out}$, which can be expressed as
$\overline{\gamma}_{r}^{out}=\operatorname{tr}\left(\boldsymbol{\rm R}_{0}\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)^{-1}\right),$ (13)
where $\boldsymbol{\rm R}_{0}=\left|\xi_{0}\right|^{2}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm F}_{\rm D}^{H}\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm A}^{H}\left(\theta_{0}\right)$. Note that the detection probability of the target increases with the output SCNR under Gaussian interference [40]. Therefore, we select the SCNR as the sensing performance metric for the ISAC system to design the hybrid beamforming. However, the hybrid beamforming matrices $\boldsymbol{\rm F}_{\rm RF}$ and $\boldsymbol{\rm F}_{\rm D}$ are highly coupled in the output SCNR, which leads to difficult handling in subsequent steps. Based on Cauchy-Schwarz inequality, we scale Eq. (13) to obtain the lower bound and define it as the input SCNR, denoted by $\overline{\gamma}_{r}^{in}$, given by
@@
where $\boldsymbol{\rm\Phi}=\left|\xi_{0}\right|^{2}\boldsymbol{\rm A}^{H}\left(\theta_{0}\right)\boldsymbol{\rm A}\left(\theta_{0}\right)$ and $\boldsymbol{\rm\Omega}=\sum_{q\in\mathcal{Q}}\left|\xi_{q}\right|^{2}\boldsymbol{\rm A}^{H}\left(\theta_{q}\right)\boldsymbol{\rm A}\left(\theta_{q}\right).$ The input SCNR $\overline{\gamma}_{r}^{in}$ more accurately depicts the quality of the sensing echo signal and is initially employed as the metric for evaluating sensing performance to design the hybrid beamforming, and then the optimal receive beamforming determined by Eq.(12) completes the echo signal filtering processing, which improves the target detection performance even further.
II-D Problem Formulation
In this work, we aim to jointly design the digital beamforming matrix $\boldsymbol{\rm F}_{\rm D}$, the analog beamforming matrix $\boldsymbol{\rm F}_{\rm RF}$, and each common information rate $C_{k},k\in\mathcal{K}$ to maximize the WSR of downlink communication users, subject to the sensing performance input SCNR, the total power budget of the system transmissions, and the unit modulus of the phase shifters. Therefore, the optimization problem of our proposed RSMA-assisted mmWave ISAC scheme can be formulated as follows:
[…]
@@ II-D Problem Formulation
where $\boldsymbol{\rm c}=\big{[}C_{1},\cdots,C_{K}\big{]}^{T}$is a vector that contains the common information rates, $\gamma_{0}$ represents the minimum SCNR threshold to satisfy target detection, and $P_{T}$ is the power budget at the ISAC-BS. Eq. (15c) ensures that each user can successfully decode the common message, and Eq. (15d) bounds the nonnegativity of all common rates. Eq. (15f) denotes the unit modulus constraint of the phase shifters.
The nonconvexity of the objective function (OF) and Eqs. (15b), (15c), and (15f) introduces a high-dimensional non-convexity to the problem P1. Moreover, the analog and digital beamforming matrices are highly coupled, and it is difficult to obtain a globally optimal solution for P1 using conventional optimization algorithms, i.e., the block coordinate descent (BCD) and the alternating optimization (AO). In the following, we seek to develop a low-complexity algorithm to solve this problem and obtain a KKT optimal solution.
III Design Algorithm
In this section, we propose the WMMSE-PDD algorithm to solve the problem P1. Specifically, we employ the WMMSE algorithm to equivalently transform the private and common information rates $R_{k}$ and $R_{c,k}$ into easy-to-handle convex function forms, and then we propose an effective joint user common rate allocation and HBF design algorithm based on the PDD method to solve the transformed problem. Finally, we summarize the algorithm and analyze its convergence and complexity.
@@ III-A Problem Reformulation
The WMMSE algorithm [42], by introducing an equalizer, can provide equivalent solutions to the original WSR maximization problems and has been widely used in the RSMA framework [32],[33],[34]. Accordingly, we apply the WMMSE algorithm to the RSMA-assisted mmWave ISAC scheme. Specifically, user $k$ first detects the common data stream $s_{c}$ by equalizer $w_{c,k}$ and then detects the private data stream $s_{k}$ by equalizer $w_{k}$ after removing the public data stream. We denote the estimates of $s_{c}$ and $s_{k}$ as $\hat{s}_{c,k}$ and $\hat{s}_{k}$, respectively, given by $\hat{s}_{c,k}=w_{c,k}y_{k}$ and $\hat{s}_{k}=w_{k}(y_{k}-\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}s_{c})$. Then, the mean square error (MSE) of estimating $s_{c}$ and $s_{k}$, denoted by $e_{c,k}$ and $e_{k}$, respectively, can be expressed as
$\hskip 20.48601pte_{c,k}$ $=\mathbb{E}\left\{\left|\hat{s}_{c,k}-s_{c}\right|^{2}\right\}$ (16)
$=\left|w_{c,k}\right|^{2}T_{c,k}-2\mathfrak{R}\left\{w_{c,k}\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}\right\}+1,$
$e_{k}$ $=\mathbb{E}\left\{\left|\hat{s}_{k}-s_{k}\right|^{2}\right\}$ (17)
$=\left|w_{k}\right|^{2}T_{k}-2\mathfrak{R}\left\{w_{k}\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},k}\right\}+1,$
where
$T_{c,k}$ $=\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}\right|^{2}+\sum_{i\in\mathcal{K}}\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}\right|^{2}+\sigma_{k}^{2},$ (18)
$T_{k}$ $=\sum_{i\in\mathcal{K}}\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}\right|^{2}+\sigma_{k}^{2}.$ (19)
The optimal equalizer can be obtained by minimizing MSE (MMSE). Letting $\frac{\partial e_{c,k}}{\partial w_{c,k}}=0$ and $\frac{\partial e_{k}}{\partial w_{k}}=0$, the optimal equalizers for $s_{c,k}$ and $s_{k}$ can be derived, denoted by $w_{c,k}^{\rm MMSE}$ and $w_{k}^{\rm MMSE}$, respectively, and expressed as $w_{c,k}^{\mathrm{MMSE}}=\boldsymbol{\rm f}_{c}^{H}\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm h}_{k}(T_{c,k})^{-1}$, $w_{k}^{\mathrm{MMSE}}=\boldsymbol{\rm f}_{{\rm D},k}^{H}\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm h}_{k}(T_{k})^{-1}$. By substituting the optimal equalizers $w_{c,k}^{\rm MMSE}$ and $w_{k}^{\rm MMSE}$ into Eqs. (16) and (17), respectively, the corresponding MMSEs, denoted by $e_{c,k}^{\rm MMSE}$ and $e_{k}^{\rm MMSE}$ can be given by
$e_{c,k}^{\rm MMSE}\triangleq\min_{w_{c,k}}e_{c,k}=\left(T_{c,k}-\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}\right|^{2}\right)(T_{c,k})^{-1},$ (20)
$e_{k}^{\rm MMSE}\triangleq\min_{w_{k}}e_{k}=\left(T_{k}-\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{\rm D,k}\right|^{2}\right)(T_{k})^{-1}.$ (21)
@@
Based on Eqs. (20) and (21), the SINRs of decoding $s_{c}$ and $s_{k}$ can be rewritten as ${\gamma}_{c,k}=1/e_{c,k}^{\rm MMSE}-1$ and ${\gamma}_{k}=1/e_{k}^{\rm MMSE}-1$. Consequently, the achievable rates of decoding $s_{c}$ and $s_{k}$ for user $k$ can be computed as $R_{c,k}=-\log_{2}\left(e_{c,k}^{\rm MMSE}\right)$ and $R_{k}=-\log_{2}\left(e_{k}^{\rm MMSE}\right)$. Next, by introducing positive weights $u_{c,k}$ and $u_{k}$, the augmented weighted MSEs (AWMSEs) of decoding $s_{c,k}$ and $s_{k}$, denoted by $\eta_{c,k}$ and $\eta_{k}$, can be expressed as $\eta_{c,k}\triangleq u_{c,k}e_{c,k}-\log_{2}\left(u_{c,k}\right)$ and $\eta_{k}\triangleq u_{k}e_{k}-\log_{2}\left(u_{k}\right)$. By substituting the optimal equalizers and weights into the AWMSEs, the rate-WMMSE relationships can be established as follows:
$\eta_{c,k}^{\rm MMSE}\triangleq\min_{u_{c,k},w_{c,k}}\eta_{c,k}=g-R_{c,k},$ (22)
$\eta_{k}^{\rm MMSE}\triangleq\min_{u_{k},w_{k}}\eta_{k}=g-R_{k},$ (23)
where $g=1/\ln 2+\log_{2}\ln 2$. The optimal equalizers for Eqs.(22) and (23) are $w_{c,k}^{\star}=w_{c,k}^{\mathrm{MMSE}}$ and $w_{k}^{\star}=w_{k}^{\mathrm{MMSE}}$, which can be derived by $\frac{\partial\eta_{c,k}}{\partial w_{c,k}}=0$ and $\frac{\partial\eta_{k}}{\partial w_{k}}=0$. And then, the optimal weights, denoted by $u_{c,k}^{\star}$ and $u_{k}^{\star}$, can be obtained by solving $\frac{\partial\eta_{c,k}(w_{c,k}^{\star})}{\partial u_{c,k}}=0$ and $\frac{\partial\eta_{k}(w_{k}^{\star})}{\partial u_{k}}=0$, which can be expressed as
$u_{c,k}^{\star}=u_{c,k}^{\rm MMSE}\triangleq(e_{c,k}^{\rm MMSE}\ln 2)^{-1},$ (24)
$u_{k}^{\star}=u_{k}^{\rm MMSE}\triangleq(e_{k}^{\rm MMSE}\ln 2)^{-1}.$ (25)
According to the rate-WMMSE relationships in Eqs.(22) and (23), the problem P1 can be equivalently transformed into the WMMSE problem, which can be expressed as follows:
[…]
@@
$\eta_{k}(\boldsymbol{\rm X})$ $=u_{k}\left|w_{k}\right|^{2}\sum_{i\in\mathcal{K}}\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm X}\boldsymbol{\rm e}_{i}\right|^{2}-2u_{k}\mathfrak{R}\left\{w_{k}\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm X}\boldsymbol{\rm e}_{k}\right\}+u_{k}\left|w_{k}\right|^{2}\sigma_{k}^{2}+u_{k}-\log_{2}\left(u_{k}\right),$ (27)
$\eta_{c,k}(\boldsymbol{\rm Z}_{k})$ $=u_{c,k}\left|w_{c,k}\right|^{2}\left(\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{c}\right|^{2}+\sum_{i\in\mathcal{K}}\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{i}\right|^{2}\right)-2u_{c,k}\mathfrak{R}\left\{w_{c,k}\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm Z}_{k}\boldsymbol{\rm e}_{c}\right\}+u_{c,k}\left|w_{c,k}\right|^{2}\sigma_{k}^{2}+u_{c,k}-\log_{2}\left(u_{c,k}\right),$ (28)
where $\boldsymbol{\rm u}=\left[u_{c,1},\cdots,u_{c,K},u_{1},\cdots,u_{K}\right]^{T}$ denotes the vector of weights and $\boldsymbol{\rm w}=\left[w_{c,1},\cdots,w_{c,K},w_{1},\cdots,w_{K}\right]^{T}$ denotes the vector of equalizers. However, there is still the nonconvex constraint in (26c) of problem P2-1, which leads to the fact that P2-1 cannot be solved directly. Regarding the relationship between the solutions of problems P1 and P2-1, the proposition is presented as follows:
Proposition 1
if $\{\boldsymbol{\rm F}_{\rm RF}^{\star},\boldsymbol{\rm F}_{\rm D}^{\star},\boldsymbol{\rm c}^{\star},\boldsymbol{\rm u}^{\star},\boldsymbol{\rm w}^{\star}\}$ is a KKT optimal solution to problem P2-1, then $\{\boldsymbol{\rm F}_{\rm RF}^{\star},\boldsymbol{\rm F}_{\rm D}^{\star},\boldsymbol{\rm c}^{\star}\}$ is a KKT optimal solution to problem P1.
[…]
In the following, we propose the PDD-based algorithm to solve problem P2-1.

### v2

@@ II-B Radar Sensing Model And Metrics
In ISAC systems, the radar sensing waveform multiplexes the communication signal $\boldsymbol{\rm x}$. We assume that the ISAC system is tasked with detecting a radar target (located at angles $\theta_{0}$) of interest in the presence of interference from $Q$ clutter located at different angles $\theta_{q},q\in\mathcal{Q}=\{1,\cdots,Q\}$. [13] Thus, the echo signal, denoted by $\boldsymbol{\rm y}_{r}$, can be expressed as
[…]
@@
where $\mathbf{a}_{t}(\theta)\in\mathbb{C}^{N_{{t}}\times 1}$ and $\mathbf{a}_{r}(\theta)\in\mathbb{C}^{N_{{r}}\times 1}$ denote the transmit and receive steering vectors, respectively. $\mathbf{z}_{r}\sim\mathcal{CN}\left(\mathbf{0}_{N_{r}},\sigma_{{z}}^{2}\mathbf{I}_{N_{r}}\right)$ represents the AWGN with zero mean and variance $\sigma_{{z}}^{2}$. $\xi_{0}$ and $\xi_{q}$ denote the complex fading coefficients of the target and the $q$-th clutter, respectively.
@@
where $\boldsymbol{\rm A}\left(\theta_{0}\right)=\sqrt{N_{r}N_{t}}\boldsymbol{\rm a}_{r}\left(\theta_{0}\right)\boldsymbol{\rm a}_{t}^{H}\left(\theta_{0}\right)$, $\boldsymbol{\rm A}\left(\theta_{q}\right)=\sqrt{N_{r}N_{t}}\boldsymbol{\rm a}_{r}\left(\theta_{q}\right)\boldsymbol{\rm a}_{t}^{H}\left(\theta_{q}\right)$. Therefore, the filtered output signal-to-clutter-plus-noise ratio (SCNR), denoted by $\gamma_{r}^{out}$, can be expressed as [14]:
$\gamma_{r}^{out}=\frac{\left|\xi_{0}\boldsymbol{\rm v}^{H}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm x}\right|^{2}}{\boldsymbol{\rm v}^{H}\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)\boldsymbol{\rm v}},$ (6)
where $\boldsymbol{\rm R}_{c}=\sum_{q\in\mathcal{Q}}\left|\xi_{q}\right|^{2}\boldsymbol{\rm A}\left(\theta_{q}\right)\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm F}_{\rm D}^{H}\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm A}^{H}\left(\theta_{q}\right)$.The receive beamforming problem of solving the output SCNR maximization is well known as the minimum variance distortionless response (MVDR) beamforming problem [15], which has a closed-form solution, denoted by $\boldsymbol{\rm v}^{\star}$, and can be expressed as $\boldsymbol{\rm v}^{\star}=\frac{\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)^{-1}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm x}}{\boldsymbol{\rm x}^{H}\boldsymbol{\rm A}^{H}\left(\theta_{0}\right)\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)^{-1}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm x}}$.
By substituting the optimal receive beamforming vector $\boldsymbol{\rm v}^{\star}$ into Eq. (6) and subsequently taking the expectation, we can derive the average output SCNR, denoted as $\overline{\gamma}_{r}^{out}$, which can be expressed as $\overline{\gamma}_{r}^{out}=\operatorname{tr}\left(\boldsymbol{\rm R}_{0}\left(\boldsymbol{\rm R}_{c}+\sigma_{z}^{2}\boldsymbol{\rm I}_{N_{r}}\right)^{-1}\right)$, where $\boldsymbol{\rm R}_{0}=\left|\xi_{0}\right|^{2}\boldsymbol{\rm A}\left(\theta_{0}\right)\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm F}_{\rm D}^{H}\boldsymbol{\rm F}_{\rm RF}^{H}\boldsymbol{\rm A}^{H}\left(\theta_{0}\right)$, and $\operatorname{tr}(\cdot)$ denote the trace of the matrix. Based on Cauchy-Schwarz inequality, we scale $\overline{\gamma}_{r}^{out}$ to obtain the lower bound and define it as the input SCNR, denoted by $\overline{\gamma}_{r}^{in}$, given by
@@
where $\boldsymbol{\rm\Phi}=\left|\xi_{0}\right|^{2}\boldsymbol{\rm A}^{H}\left(\theta_{0}\right)\boldsymbol{\rm A}\left(\theta_{0}\right)$ and $\boldsymbol{\rm\Omega}=\sum_{q\in\mathcal{Q}}\left|\xi_{q}\right|^{2}\boldsymbol{\rm A}^{H}\left(\theta_{q}\right)\boldsymbol{\rm A}\left(\theta_{q}\right).$ The input SCNR $\overline{\gamma}_{r}^{in}$ improves the target detection performance even further.
II-C Problem Formulation
In this work, we aim to jointly design the HBF matrix and each common information rate $C_{k},k\in\mathcal{K}$ to maximize the WSR. The optimization problem can be formulated as follows:
$\textbf{{{P}}1:}\max_{\boldsymbol{\rm F}_{\rm RF},\boldsymbol{\rm F}_{\rm D},\boldsymbol{\rm c}}$ $\sum_{k\in\mathcal{K}}\alpha_{k}(C_{k}+R_{k})$ (8a)
@@ II-C Problem Formulation
where $\left\|\cdot\right\|_{p}$ denote $l_{p}$ norm operation, $\boldsymbol{\rm c}=\big{[}C_{1},\cdots,C_{K}\big{]}^{T}$ is a vector that contains the common information rates, $\gamma_{0}$ represents the minimum SCNR threshold to satisfy target detection, and $P_{T}$ is the power budget at the ISAC-BS. Eq. (8c) ensures that each user can successfully decode the common message, and Eq. (8d) bounds the nonnegativity of all common rates. Eq. (8f) denotes the unit modulus constraint of the phase shifters.
The nonconvexity of the objective function (OF) complicates finding a global optimum for P1 using conventional optimization algorithms like the block coordinate descent (BCD) and the alternating optimization (AO).
III Proposed WMMSE-PDD Algorithm
In this section, we propose the WMMSE-PDD algorithm to solve the problem P1.
@@ III-A Problem Reformulation
According to [10], the minimizing mean square error (MMSE) of estimating $s_{c}$ and $s_{k}$, denoted by $e_{c,k}^{\rm MMSE}$ and $e_{k}^{\rm MMSE}$ can be given by
$\begin{cases}&\hskip-2.84544pte_{c,k}^{\rm MMSE}\!=\!(T_{c,k}\!-\!\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}\right|^{2})(T_{c,k})^{-1},\\ &\hskip-2.84544pte_{k}^{\rm MMSE}\!=\!(T_{k}\!-\!\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{\rm D,k}\right|^{2})(T_{k})^{-1}.\end{cases}$ (9)
@@
where $T_{c,k}=\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}\right|^{2}+\sum_{i\in\mathcal{K}}\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}\right|^{2}+\sigma_{k}^{2}$ and $T_{k}=\sum_{i\in\mathcal{K}}\left|\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}\right|^{2}+\sigma_{k}^{2}$. Then, we introduce positive weights $u_{c,k}$ and $u_{k}$, the augmented weighted MSEs (AWMSEs) of decoding $s_{c,k}$ and $s_{k}$, denoted by $\eta_{c,k}$ and $\eta_{k}$. By substituting the optimal equalizers and weights into the AWMSEs, we can obtain $\eta_{c,k}^{\rm MMSE}\triangleq\min\limits_{u_{c,k},w_{c,k}}\eta_{c,k}=g-R_{c,k}$ and $\eta_{k}^{\rm MMSE}\triangleq\min\limits_{u_{k},w_{k}}\eta_{k}=g-R_{k}$, where $g=1/\ln 2+\log_{2}\ln 2$. Based on this, the problem P1 can be equivalently transformed into the WMMSE problem, which can be expressed as follows:
$\textbf{{{P}}2-1:}\min_{\begin{subarray}{c}\boldsymbol{\rm F}_{\rm RF},\boldsymbol{\rm F}_{\rm D},\boldsymbol{\rm c},\\ \boldsymbol{\rm u},\boldsymbol{\rm w}\end{subarray}}$ $\sum_{k\in\mathcal{K}}\alpha_{k}(\eta_{k}-C_{k})$ (10a)
@@
where $\boldsymbol{\rm u}=\left[u_{c,1},\cdots,u_{c,K},u_{1},\cdots,u_{K}\right]^{T}$ denotes the vector of weights and $\boldsymbol{\rm w}=\left[w_{c,1},\cdots,w_{c,K},w_{1},\cdots,w_{K}\right]^{T}$ denotes the vector of equalizers. However, there is still the nonconvex constraint in (10c) of problem P2-1, which leads to the fact that P2-1 cannot be solved directly.